Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 600,000 shares of common stock were outstanding. The interest rate on the bond payable was 10%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of this year was $26. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,130 $ 1,310
Accounts receivable, net 9,600 8,100
Inventory 12,100 11,200
Prepaid expenses 780 660
Total current assets 23,610 21,270
Property and equipment:
Land 11,000 11,000
Buildings and equipment, net 50,014 39,512
Total property and equipment 61,014 50,512
Total assets $ 84,624 $ 71,782
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 18,600 $ 17,900
Accrued liabilities 1,050 820
Notes payable, short term 0 190
Total current liabilities 19,650 18,910
Long-term liabilities:
Bonds payable 9,300 9,300
Total liabilities 28,950 28,210
Stockholders' equity:
Common stock 2,000 2,000
Additional paid-in capital 4,000 4,000
Total paid-in capital 6,000 6,000
Retained earnings 49,674 37,572
Total stockholders' equity 55,674 43,572
Total liabilities and stockholders' equity $ 84,624 $ 71,782
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 73,000 $ 64,000
Cost of goods sold 34,000 34,000
Gross margin 39,000 30,000
Selling and administrative expenses:
Selling expenses 11,000 10,200
Administrative expenses 6,500 6,900
Total selling and administrative expenses 17,500 17,100
Net operating income 21,500 12,900
Interest expense 930 930
Net income before taxes 20,570 11,970
Income taxes 8,228 4,788
Net income 12,342 7,182
Dividends to common stockholders 240 450
Net income added to retained earnings 12,102 6,732
Beginning retained earnings 37,572 30,840
Ending retained earnings $ 49,674 $ 37,572

Required:

Compute the following financial ratios for this year:

1. Times interest earned ratio.

2. Debt-to-equity ratio.

3. Equity multiplier.

(For all requirements, round your answers to 2 decimal places.)
Answer #1
1. Times interest earned ratio = Net operating income / Interest expense $21,500 / 930 = $23.12
2. Debt to equity ratio = Total liabilities / Total equity $28,950 / 55,674 = 0.52
3.

Equity multiplier = Average assets / Average equity

Average = Beginning balance+Ending balance / 2

$78,203 / 49,623 = 1.58